# Scalable and shallow quantum circuits encoding probability distributions informed by asymptotic entanglement analysis

### v1

@@ 1 Introduction
Figure 1: (a) A real or complex function $f(x)$ on $[0,1]$ is discretized on a binary grid $x=\sum_{i}\sigma_{i}2^{-i}$. Successive digits $\sigma_{i}$ of the binary expansion describe decreasing spatial scales: parts of the interval $[0,1]$ sharing the same value of $\sigma_{i}$ are shown in green and blue. (b) Discretized function can be represented as a rank-$N$ tensor $f_{\sigma_{1}\sigma_{2}\dots\sigma_{N}}$. (c) A matrix product state (MPS) can approximate $f_{\sigma_{1}\sigma_{2}\dots\sigma_{N}}$, but the bond dimensions $l_{k}$, and thus the entanglement across bond $k$, may be exponential in general case. We show that for smooth functions the entanglement decays exponentially with $k$, effectively decoupling fine spatial scales. For the very largest scales the behavior is non-universal and depends on the localization properties of the function. (d) We use these analytical results to construct optimized quantum circuits with $O(N)$ gates generating the state encoding the function $f$. The algorithm minimizes the number of 2-qubit gates based on the entanglement present at different scales.
@@ 3.1 General results
In Appendix A.4 we further observe that if the function is not smooth, but only continuous up to $r$-th derivative (i.e. in class $C^{r}$), then our previous results hold for entanglement spectra up to order $r+1$, while the $r+2$ and higher orders are non-universal.
@@
Figure 3: Entanglement decay for exponential and polynomial localization: a) The entanglement spectra $\Lambda_{k,1}$ as a function of the bond position $k$ for the exponentially localized one-sided normal ($\mu=0$, $\sigma=1$) distribution truncated to the interval $[0,16]$, for $N=18$ qubits. Note the super-exponential decay of entanglement for small $k$ (and the universal, exponential, smoothness-controlled regime at large $k$). b) The entanglement spectra $\Lambda_{k,1}$ as a function of the bond position $k$ for the Lévy distribution (with $c=1$ and interval $[0,2^{19}]$), on $N=27$ qubits. For small $k$ the entanglement decays super-exponentially, but due to only polynomial localization slower than for normal distribution, delaying the onset of the universal regime at large $k$.
@@ 3.3 Encoding dependence on the localization and support of the function
Figure 5: The dependence of the infidelity on the number and position of the origin of the V-layers in the encoding quantum circuits on an ideal simulator. Here we show data for the Lévy distribution with $c=1$ on support truncated to $L=2^{10}$ using $N=18$ qubits. The accuracy increases with additional disentangling layers (but see Fig.12 for the effects of noise), and significantly depends on the origin of the V-layers. The effect of the latter, which also reduces to overall circuit depth, has effect comparable to adding a new disentangling layer.
@@
Figure 6: (a) The log-log plot of infidelity of a 1-layer quantum circuit as a function of the size of the truncated support interval $L$. We fix the discretization step size, with the qubit number given by $N=10+\lceil\log_{2}L\rceil$. Here we consider the truncated normal distribution with $\sigma=2$, show the numerically computed infidelity, and compare it to the analytical estimates. The label “Analytical” refers to Eq.(17) and “Analytical with correction” to prediction taking into account localization properties (see the main text). Note the very low infidelity obtained already for this shallow circuit. (b) The numerical infidelity of the 1-layer quantum circuit for the truncated Lévy distribution with parameter $c=2^{5}$ as a function of the support interval $L$. The discretization step size is fixed, with the qubit number given by $N=10+\lceil\log_{2}L\rceil$.
@@ 4 Constructing the encoding circuit using the MPS representation
Figure 7: The encoded Lévy distribution ($c=1$) for different truncated support and discretization (number of qubits) sizes on an ideal noiseless simulator. The amplitudes of the quantum state created using a 2-layer encoding circuit are compared to the exact distribution.
@@
Figure 8: Encoding circuits execution on the ibm_torino QPU. (a) Lévy distribution with $c=50$ truncated to $L=2^{9}$ with $N=20$ qubits (b) Lévy distribution with $c=1310$ on $L=2^{14}$ with $N=25$ qubits (c) Log-normal distribution with on $N=25$ qubits (d) Gamma distribution with shape parameter $k=1$ on $N=25$ qubits. In each case 5000 shots were used. All executions pass the KS test with 200 samples. The better of the 1- and 2-layer circuits used is shown.
@@ 5 Numerical tests of the encoding circuits on simulators and QPUs
In Fig.5 we examine the influence of using multiple disentangling V-layers in the quantum circuit construction, and the sensitivity to the V-layer origin position, on the example of the Lévy dsitribution. We observe that the fidelity can be systematically improved by adding additional disentangling layers (note, though, that in the presence of noise this is not true, as the increased circuit depth may result in a lower fidelity in practice). Furthermore, for more than a single layer we also see a strong dependence on the layer origin, as expected from the entanglement analysis in Sec.3.1: the effect is comparable in size to adding another disentangling layer, and hence the freedom to chose it is very important in practice. Note also that the minimal circuit depth is achieved with the origin on the middle qubits, further emphasizing the usefulness of this strategy.
@@
Figure 9: Tensor Cross Interpolation (TCI) constructed 2-layer encoding circuits for the Gamma and Lévy distributions on 50 qubits, executed on the ibm_marrakesh QPU. We used in both cases $L=10^{11}$, the scale factor $c=5\times 10^{9}$, and 5000 shots on the QPU. While these very large circuits no longer pass the quantitative statistical tests, the qualitative features of the distributions are still reproduced.
@@ B.2 Exact functional coefficients for probability distributions
$g_{1}(\sqrt{p(x)_{normal}})=\frac{L^{2}}{4\sigma^{2}}\left[1-\sqrt{\frac{L^{2}}{2\sigma^{2}}}\exp{\left(-\frac{L^{2}}{8\sigma^{2}}\right)}/\left(\Gamma(1/2)-\Gamma(1/2,L^{2}/8\sigma^{2})\right)\right]$ (49)
$g_{1}(\sqrt{p(x)_{log-normal}})=\frac{\exp{\left(-2\mu+2\sigma^{2}\right)}}{4\sigma^{2}\left[\sqrt{2\pi}-\Gamma(1/2,(\log{(L\exp(-\mu))}/\sigma)^{2}/2)\right]}\times\\ \times[(1+\sigma^{2})(\sqrt{2\pi}-\Gamma(1/2,(\log{(L\exp(-\mu))}/\sigma+2\sigma)^{2}/2))-\\ -\left(\frac{\log{(L\exp{(-\mu)})}}{\sigma}\right)\exp{[-(\log{(L\exp{[-\mu]})}/\sigma+\sigma)^{2}/2]}]-\\ -\exp{\{-2\mu+\sigma^{2}-(\log{(L\exp(-\mu))}/\sigma+\sigma)^{2}\}}/(4\sigma^{2}\times\left[\sqrt{2\pi}-\Gamma(1/2,(\log{(L\exp(-\mu))}/\sigma)^{2}/2)\right]^{2})$ (50)
@@
For the Lévy distribution we computed $g_{1}(f)$ for the discretization on the finite interval $L$. The leading term $3L^{2}/c^{2}$ defines the characteristic scale $k=\log{L/c}$, where the entanglement spectra become small. Other terms define the subleading corrections due to the finiteness of the interval. Here $\Gamma(s,x)$ is the incomplete $\Gamma$ function, which comes from the normalization of the Lévy distribution on the finite interval.
@@
$g_{2}(\sqrt{p(x)_{normal}})=\frac{L^{4}}{8\sigma^{4}}-\frac{L^{4}}{4\sigma^{4}}\left(\sqrt{\frac{L^{2}}{32\sigma^{2}}}+\sqrt{\frac{L^{2}}{8\sigma^{2}}}^{3}\right)\exp{\left[-\frac{L^{2}}{8\sigma^{2}}\right]}/(\Gamma(1/2)-\Gamma(1/2,L^{2}/8\sigma^{2}))-\\ -\frac{L^{6}}{32\sigma^{6}}\exp{\left[-\frac{L^{2}}{4\sigma^{2}}\right]}(\Gamma(1/2)-\Gamma(1/2,L^{2}/8\sigma^{2}))^{2}$ (52)
@@ Appendix C Entanglement and function localization
While we do not provide here rigorous results, it is possible to characterize the large scale entanglement slightly better for specific classes of functions. An important case are the localized functions. Informally a function is localized exponentially or polynomially if the integral of the function over increasing subsets containing the domain of localization approaches unity (we assumed normalized function) with the remainder decaying polynomially or exponentially in the size of the subsets.
@@ Appendix D Tensor cross interpolation
Figure 11: Tensor Cross Interpolation (TCI) constructed 2-layer encoding circuits for the Lévy distributions on 50 and 64 qubits, executed on the ibm_marrakesh QPU. We used in both cases $L=10^{11}$, the scale factor $c=5\times 10^{9}$, and 5000 shots on the QPU. While these very large circuits no longer pass the quantitative statistical tests, the qualitative features of the distributions are still reproduced.

### v2

@@ 1 Introduction
Figure 1: a) A real or complex function $f(x)$ on $[0,1]$ is discretized on a binary grid $x=\sum_{i}\sigma_{i}2^{-i}$. Successive digits $\sigma_{i}$ of the binary expansion describe decreasing spatial scales: parts of the interval $[0,1]$ sharing the same value of $\sigma_{i}$ are shown in green and blue. b) Discretized function can be represented as a rank-$N$ tensor $f_{\sigma_{1}\sigma_{2}\dots\sigma_{N}}$. c) A matrix product state (MPS) can approximate $f_{\sigma_{1}\sigma_{2}\dots\sigma_{N}}$, but the bond dimensions $l_{k}$, and thus the entanglement across bond $k$, may be exponential in general case. We show that for smooth functions the entanglement decays exponentially with $k$, effectively decoupling fine spatial scales. For the very largest scales the behavior is non-universal and depends on the localization properties of the function. d) We use these analytical results to construct optimized quantum circuits with $O(N)$ gates generating the state encoding the function $f$. The algorithm minimizes the number of 2-qubit gates based on the entanglement present at different scales.
@@ 3.1 General results
In Appendix A.4 we further observe that if the function is not smooth, but has absolutely continuous derivatives up to $r$-th order, then our previous results hold for entanglement spectra up to order $r+1$, while the $r+2$ and higher orders are non-universal.
@@
Figure 3: Entanglement decay for exponential and polynomial localization: a) The entanglement spectra $\Lambda_{k,1}$ as a function of the bond position $k$ (lines are for guidance only) for the exponentially localized one-sided normal ($\mu=0$, $\sigma=1$) distribution truncated to the interval $[0,16]$, for $N=18$ qubits. Note the super-exponential decay of entanglement for small $k$ (and the universal, exponential, smoothness-controlled regime at large $k$). b) The entanglement spectra $\Lambda_{k,1}$ as a function of the bond position $k$ for the Lévy distribution (with $c=1$ and interval $[0,2^{19}]$), on $N=27$ qubits. For small $k$ the entanglement decays super-exponentially, but due to only polynomial localization slower than for normal distribution, delaying the onset of the universal regime at large $k$.
@@ 3.3 Encoding dependence on the localization and support of the function
Figure 5: The dependence of the infidelity on the number and position of the origin of the V-layers in the encoding quantum circuits on an ideal simulator (lines between points for guidance only). Here we show data for the Lévy distribution with $c=1$ on support truncated to $L=2^{10}$ using $N=18$ qubits. The accuracy increases with additional disentangling layers (but see Fig.12 for the effects of noise), and significantly depends on the origin of the V-layers. The effect of the latter, which also reduces to overall circuit depth, has effect comparable to adding a new disentangling layer.
@@
Figure 6: a) The log-log plot of infidelity of a 1-layer quantum circuit as a function of the size of the truncated support interval $L$. We fix the discretization step size, with the qubit number given by $N=10+\lceil\log_{2}L\rceil$. Here we consider the truncated normal distribution with $\sigma=2$, show the numerically computed infidelity, and compare it to the analytical estimates. The label “Analytical” refers to Eq.(17) and “Analytical with correction” to prediction taking into account localization properties (see the main text). Note the very low infidelity obtained already for this shallow circuit. b) The numerical infidelity of the 1-layer quantum circuit for the truncated Lévy distribution with parameter $c=2^{5}$ as a function of the support interval $L$. The discretization step size is fixed, with the qubit number given by $N=10+\lceil\log_{2}L\rceil$.
@@ 4 Constructing the encoding circuit using the MPS representation
Figure 7: The encoded Lévy distribution ($c=1$) for different truncated support and discretization (number of qubits) sizes on an ideal noiseless simulator. Amplitudes of the quantum state created using a 2-layer encoding circuit are compared to the exact distribution. In a) for $N=6$ qubits and $L=1$; in b) for $N=10$ qubits and $L=5$.
@@
Figure 8: Encoding circuits execution on the ibm_torino QPU. a) Lévy distribution with $c=50$ truncated to $L=2^{9}$ with $N=20$ qubits b) Lévy distribution with $c=1310$ on $L=2^{14}$ with $N=25$ qubits c) Log-normal distribution with on $N=25$ qubits d) Gamma distribution with shape parameter $k=1$ on $N=25$ qubits. In each case 5000 shots were used. All executions pass the KS test with 200 samples. The better of the 1- and 2-layer circuits used is shown.
@@ 5 Numerical tests of the encoding circuits on simulators and QPUs
In Fig.5 we examine the influence of using multiple disentangling V-layers in the quantum circuit construction, and the sensitivity to the V-layer origin position, on the example of the Lévy distribution. We observe that the fidelity can be systematically improved by adding additional disentangling layers (note, though, that in the presence of noise this is not true, as the increased circuit depth may result in a lower fidelity in practice). Furthermore, for more than a single layer we also see a strong dependence on the layer origin, as expected from the entanglement analysis in Sec.3.1: the effect is comparable in size to adding another disentangling layer, and hence the freedom to chose it is very important in practice. Note also that the minimal circuit depth is achieved with the origin on the middle qubits, further emphasizing the usefulness of this strategy.
@@
Figure 9: Tensor Cross Interpolation (TCI) constructed 2-layer encoding circuits for the a) Gamma and b) Lévy distributions on 50 qubits, executed on the ibm_marrakesh QPU. We used in both cases $L=10^{11}$, the scale factor $c=5\times 10^{9}$, and 5000 shots on the QPU. While these very large circuits no longer pass the quantitative statistical tests, the qualitative features of the distributions are still reproduced.
@@ B.2 Exact functional coefficients for probability distributions
$g_{1}(\sqrt{p(x)_{normal}})=\frac{L^{2}}{4\sigma^{2}}\left[1-\sqrt{\frac{L^{2}}{2\sigma^{2}}}\frac{\exp{\left(-\frac{L^{2}}{8\sigma^{2}}\right)}}{\left(\Gamma(1/2)-\Gamma(1/2,L^{2}/8\sigma^{2})\right)}\right]$ (49)
$g_{1}(\sqrt{p(x)_{log-normal}})=\frac{L^{2}\exp{\left(-2\mu+2\sigma^{2}\right)}}{4\sigma^{2}\left[\sqrt{2\pi}-\Gamma\Bigl{(}1/2,\frac{(\log{(L\exp(-\mu))}/\sigma)^{2}}{2}\Bigr{)}/\sqrt{2}\right]}\times\\ \times\biggl{[}(1+\sigma^{2})\biggl{(}\sqrt{2\pi}-\Gamma\biggl{(}1/2,\frac{(\log{(L\exp(-\mu))}/\sigma+2\sigma)^{2}}{2}\biggr{)}/\sqrt{2}\biggr{)}+\\ +\left((2-\sqrt{2})\sigma-\frac{\log{(L\exp{(-\mu)})}}{\sqrt{2}\sigma}\right)\exp{[-(\log{(L\exp{[-\mu]})}/\sigma+2\sigma)^{2}/2]}\biggr{]}-\\ -\frac{L^{2}\exp{\{-2\mu+\sigma^{2}-(\log{(L\exp(-\mu))}/\sigma+\sigma)^{2}\}}}{4\sigma^{2}\times\left[\sqrt{2\pi}-\Gamma\Bigl{(}1/2,\frac{(\log{(L\exp(-\mu))}/\sigma)^{2}}{2}\Bigr{)}/\sqrt{2}\right]^{2}}$ (50)
@@
For the Lévy distribution we computed $g_{1}(f)$ for the discretization on the finite interval $L$. The leading term $21L^{2}/8c^{2}$ defines the characteristic scale $k=\log{L/c}$, where the entanglement spectra become small. Other terms define the subleading corrections due to the finiteness of the interval. Here $\Gamma(s,x)$ is the incomplete $\Gamma$ function, which comes from the normalization of the Lévy distribution on the finite interval.
@@
$g_{2}(\sqrt{p(x)_{normal}})=\frac{L^{4}}{8\sigma^{4}}-\frac{L^{4}}{4\sigma^{4}}\left(\sqrt{\frac{L^{2}}{32\sigma^{2}}}+\sqrt{\frac{L^{2}}{8\sigma^{2}}}^{3}\right)\frac{\exp{\left[-\frac{L^{2}}{8\sigma^{2}}\right]}}{(\Gamma(1/2)-\Gamma(1/2,L^{2}/8\sigma^{2}))}-\\ -\frac{L^{6}}{32\sigma^{6}}\frac{\exp{\left[-\frac{L^{2}}{4\sigma^{2}}\right]}}{(\Gamma(1/2)-\Gamma(1/2,L^{2}/8\sigma^{2}))^{2}}$ (52)
@@ Appendix C Entanglement and function localization
While we do not provide here rigorous results, it is possible to characterize the large scale entanglement slightly better for specific classes of functions. An important case are the localized functions. Informally, a function is localized exponentially or polynomially if the integral of the function over increasing subsets containing the domain of localization approaches unity (we assumed normalized function) with the remainder decaying polynomially or exponentially in the size of the subsets.
@@ Appendix D Tensor cross interpolation
Figure 11: Tensor Cross Interpolation (TCI) constructed 2-layer encoding circuits for the Lévy distributions on 50 (left) and 64 (right) qubits, executed on the ibm_marrakesh QPU. We used in both cases $L=10^{11}$, the scale factor $c=5\times 10^{9}$, and 5000 shots on the QPU. While these very large circuits no longer pass the quantitative statistical tests, the qualitative features of the distributions are still reproduced.